graph and check to solve the linear system. 3x+y=5 -x+y=-7
Answer by mananth(16949)

graph and check to solve the linear system. 3x+y=5 -x+y=-7
...
3x+5y=5................1
-x+y=-7................2
...
multiply (2) by 3
-3x+3y=-21
add to (1)
8y=-16
y=-2
..
plug value of y in 2
-x-2=-7
-x=-5
x=5
...